OpenStudy (anonymous):

Write the quadratic function with the following transformations: vertical shrink by 3, left 4, down 9

OpenStudy (anonymous):

@ganeshie8

OpenStudy (lucaz):

what is the quadratic function? quadratic in the form y=(x-h)^2+k shrink by 3 = multiply (x-h)^2 by 1/3 left 4 = (x-h+4)^2 down 9 = k-9